OpenStudy (anonymous):

Please help me! Will give medal and fan! =) What is the equation, in standard form, of a parabola that contains the following points? (–2, -20), (0, -4), (4, -20)

OpenStudy (anonymous):

y = 3x^2 – 2x + 2

OpenStudy (anonymous):

idk why those boxes show up but its suppose to be a minus sign there...

OpenStudy (anonymous):

How do I solve it though? I have to show work and I would like to understand it too...

OpenStudy (bohotness):

-1408

OpenStudy (anonymous):

the general form of a parabola is \[y=ax^2+bx+c\] Now we can substitue the points x = -2, y = -20 \[-20=4a-2b+c\] x =0, y = -4 \[-4=c\] x = 4, y = -20 \[-20=16a+4b+c\] We can equate the first and last equations to each other \(4a-2b+c=16a+4b+c\), that is \(4a-2b=16a+4b\), or \(2a-b=8a+2b\), which amounts to \(6a=-3b\), or \(2a=-b\) Tis can be backsubstituted in the first and last equations, together with the found value of \(-4=c\) \(-20=-2b-2b-4\), or \(4b=16\), so that \(b=4\) therefore, \(a=-2\), since we had \(2a=-b\) This can be checked by substituting all found values into the third eqation \(-20 = 16 times {-2} - 4 times 4 - 4\), or \(-20 = -32+16-4\) So the equation for the parabola is \[y=-2x^2+4x-4\] https://s3.amazonaws.com/grapher/exports/cf3ghpqsxk.png